# KLAN: Kuaishou Landing-page Adaptive Navigator

### v1

@@ Abstract.
Personalized Landing Page, Uplift Model, Reinforment Learning
ccs: Information systems Recommender systems
@@ 1. Introduction
With the prevalence of Web 2.0 and mobile devices, a massive number of users have joined online platforms such as Amazon, TikTok, and Kuaishou for shopping, content creation, and social networking. On these platforms, multiple functional and channel pages are configured to cater to users’ diverse needs. For example, Kuaishou offers various types of pages including the ”Featured” page that presents recent and personalized recommendations to users and the ”Explore” page that provides less personalized but more diverse video recommendations. Inside each of these pages, users interact with the recommender system with consecutive requests, receiving sequential recommendations and providing feedback. In this sense, we consider the multi-page online recommender system as a two-stage interaction paradigm between the user and the platform (shown in Figure 1): an initial page navigation stage that navigates users to a specific functional or channel page, and a subsequent in-page interaction stage where users sequentially engage with customized content on the specific page.
@@
While most existing approaches for video recommendation platforms focus on the sequential recommendation task in the second in-page interaction stage (Kang and McAuley, 2018; Sun et al., 2019; Mao et al., 2025, 2024; Guo et al., 2023; Meng et al., 2023a, b, 2024; Chang et al., 2023; Chen et al., 2021; Pi et al., 2020), there has been little investigation on how to achieve better page navigation in the first stage. However, the initial page navigation stage significantly impacts the user experience in the system. Firstly, it determines whether the following interactions can occur and directly influences the user’s first impression of the online service, which is reflected by short-term metrics like the Page Drop-off Ratio (PDR), as well as the long-term metrics such as Daily Active Users (DAU) and LifeTime (LT). Here, we define the Page Drop-off Ratio (PDR) as the probability that a user immediately leaves the assigned initial page upon app entry. Furthermore, users may typically enter the app multiple times in one day (Figure 2-c), so effective page navigation may choose different initial pages that accurately address the users’ dynamic demands upon each entry. Yet, inaccurate navigation could cause the users to frequently switch pages (i.e. the page drop-off), so it is also critical to avoid this negative impact on the user experience. Additionally, empirical evidence from Kuaishou, as shown in Figure 2-a and 2-b, demonstrates that approximately 42% of users engage across multiple pages and consistently outperform single-page users in terms of DAU and LT. While the remaining 58% of users exhibit single-page activities within a single day (shown in Figure 2-b), we believe that when the system properly introduces new pages to these users, it is possible to stimulate their potential in multi-page consumption, which improves the exposure of diverse video contents on different pages and eventually motivates the sustainable advancement of the platform’s ecosystem.
@@
  • Inter-day static page preferences modeling: User’s static page preference expresses the long-term stable preferences towards each pages, which is a critical feature to model the landing page of a user’s first entry on each day. Empirically, users have imbalanced interactions across pages, which introduces considerable exposure bias. For example, we observe that the vast majority of samples are available on the ”Featured” page. On the other hand, different pages are heterogeneous in presentation format, including content and interaction patterns, which results in vastly different user response distributions across pages (Section A.1). Given such heterogeneity, modeling user engagement for each page in isolation fails to capture the underlying causal relationships. Instead, a holistic perspective is needed—one that explicitly accounts for how assigning a user to one page over another influences their subsequent behavior. While uplift modeling provides a principled framework for estimating such treatment effects, existing approaches often rely on joint optimization over multiple objectives, which introduces gradient conflict problem (detailed in Appendix A.3), ultimately leading to suboptimal performance in practice.
@@ 2.1. Uplift Model for Video Recommendation Platform
Uplift models aim to identify the target user groups for each specific treatment by estimating individual treatment effects (ITE). Existing approaches can be categorized into three main methodologies: Meta-learner based methods, Tree-based methods and Neural Networks based methods. Meta-learner based methods (Künzel et al., 2019; Yao et al., 2021) constitute a classical framework for uplift modeling. The S-learner (Künzel et al., 2019) treats the treatment variable as a feature within a single model, while the T-learner (Künzel et al., 2019) builds separate models for treatment and control groups. Tree-based methods (Athey and Imbens, 2016; Nandy et al., 2023; Radcliffe and Surry, 2011; Wang et al., 2015) utilize specific tree or forest structures, setting direct computational methods for incremental uplift modeling. Due to the ability to capture complex treatment-response relationships and estimate more accurate ITE, neural networks have emerged as a popular framework for uplift modeling. Existing Neural Network-based uplift methods (Bica et al., 2020; Louizos et al., 2017; Shi et al., 2019; Wei et al., 2024; Zhou et al., 2023) can be categorized into two categories: joint modeling of treatment and modeling of specific treatment causal effects. The former estimates the non-treatment score through features and incorporates the treatment feature into the treatment modeling based on this score. For instance, EUEN (Ke et al., 2021) uses different sub-neural networks to estimate the conversion probability of the control group and the uplift effect respectively. The latter focuses on modeling the causal effects of specific treatment. DESCN (Zhong et al., 2022) designs an end-to-end deep cross network to comprehensively capture the information of intervention propensity, response and hidden intervention effects through multi-task learning. CDUM (Meng et al., 2024) expands the treatment representation for guidance and indicator within a multi-task framework and proposes an online real-time feature learning network based on video recommendation scenarios. However, existing methods either overemphasize the collaborative effects or the causal effects, making it difficult to adapt to short video recommendation scenarios which involves heterogeneous multi-treatment (Appendix A.3).
@@ 2.2. Reinforcement Learning for Video Recommendation Platform
Unlike traditional methods optimizing for the one-step recommendation, reinforcement learning (Sutton and Barto, 1998) approaches model video recommendation as a Markov Decision Process, optimizing cumulative rewards to maximize long-term user satisfaction. Reinforcement learning (RL) approaches can be broadly categorized into on-policy and off-policy methods. Due to the high policy variance and low sample efficiency of on-policy methods, off-policy methods have emerged as the predominant choice in industrial application. Among off-policy methods, policy gradient approaches (Chen et al., 2019b, a; Ge et al., 2021, 2022; Li et al., 2022; Xian et al., 2019) suffer from the mismatch between the behavior policy and the target policy, leading to substantial discrepancies between offline evaluation and online performance. Value gradient approaches (van Hasselt et al., 2015; Pei et al., 2019; Taghipour et al., 2007; Zhao et al., 2021, 2018b), prominently represented by DQN (van Hasselt et al., 2015), have demonstrated superior stability and data efficiency by leveraging experience replay. However, the traditional DQN strrugle with the overestimation bias due to the exposure imbalance in our proposed PLPM scenario. To address it, conservative reinforcement learning algorithms such as CQL (Kumar et al., 2020) and IQL (Kostrikov et al., 2021) have been proposed, incorporating pessimistic value estimation to handle distributional shift. Despite the theoretical advantages of CQL, its practical deployment reveals a critical hyperparameter sensitivity issue. Specifically, an excessively large conservative coefficient leads to overly pessimistic policies that fail to stimulate multi-page user engagement, while an insufficient coefficient approximates standard DQN behavior, resulting in policy instability that violates online serving constraints and degrades user experience. Furthermore, our empirical analysis on Kuaishou reveals that intra-day RL decision exhibit a ”tidal phenomenon” (detailed in Appendix A.4). Motivated by this, we introduce a Dynamic Conservative Coefficient, leveraging intra-day real-time features strongly correlated with the ”tidal phenomenon” (such as request hour and daily app entry frequency) to automatically balance the exploration-exploitation trade-off.
@@ 3. The PLPM Problem Formulation
We define the three sub-tasks of PLPM problem as inter-day modeling of user’s static page preferences, the intra-day modeling of user’s dynamic page interest transition and the adaptive modulation of both components.
@@ 3.2. The Intra-day Modeling Task
  • $\mathcal{S}$: The continuous representation space of the user state. Each state $s_{t}$ encodes the user’s real-time features and daily features upon app entry at step $t$.
@@
The goal of this MDP is to learn an optimal policy $\pi^{*}$:$\mathcal{S}\rightarrow\mathcal{A}$ that maximizes the expected discounted return:
@@ 4.1.2. Multi-treatment Joint Optimization
Through Treatment-Specific Uplift Model, we obtain the estimated output pairs $[\hat{y}_{i}^{k},\hat{y}_{i}^{0,k}]$ and the latent representation pairs $[\mathbf{z}_{i}^{k},\mathbf{z}_{i}^{0,k}]$ from each branch. For training, we adopt the mean square error as the main training loss. To improve the robustness and effectiveness of ISP, we apply KL divergence to constrain non-treatment representations across branches. Therefore, the total training loss for ISP is as follows:
(5) =$\mathcal{L}=\begin{cases}\frac{1}{N}\sum_{i=1}^{N}\sum_{k^{\prime}=1}^{K}\left(\left(y_{i}^{k}-\hat{y}_{i}^{0,k^{\prime}}\right)^{2}+D_{\mathrm{KL}}\left(z_{i}^{0,k^{\prime}}\,\middle\|\,{z}_{i}^{0,*}\right)\right),&k=0\\ \frac{1}{N}\sum_{i=1}^{N}\left(y_{i}^{k}-\hat{y}_{i}^{k}\right)^{2},&k\neq 0\end{cases}$
@@ 4.1.3. Prediction
During inference, we iteratively replace the treatment feature $\mathbf{t}i$ to obtain predictions for all treatments: $\{\hat{y}_{i}^{k^{\prime}},\hat{y}_{i}^{0,k^{\prime}}\}_{k^{\prime}=0}^{K}$.
@@ 5.1. Dataset
To model users’ dynamic interest transitions in personalized landing page scenarios, we construct an hourly-level dataset tailored for reinforcement learning (RL) tasks. This dataset is used to train the KLAN-IIT module, which captures temporal patterns in user behavior across app entry sessions. Each instance corresponds to a complete ”app entry session” (defined as the sequence of interactions that starts with the display of a landing page when the user enters the app and ends when the user exits) and the required features of the corresponding next session. The dataset is organized in the standard RL format $(s_{t},a_{t},r(s_{t},a_{t}),s_{t+1})$, where $a_{t}$ denotes the landing page delivered at the beginning of the session. The state $s_{t}$ includes both real-time contextual features—such as current-day page-level consumption statistics, the number of live broadcasts by followed creators at the time of entry, and the number of prior app entries on the same day—and historical statistical features, including aggregated page consumption behavior over the past $n$ days. The reward $r(s_{t},a_{t})$ is defined as the app usage duration of the current session, serving as a proxy for user engagement. This dataset is constructed from hourly online logs collected in a real-world industrial environment, ensuring high temporal resolution and behavioral fidelity.
@@
Figure 4. Illustration of the online pipeline.
5.2. Online A/B Test
5.2.1. System Description.
5.2.2. Scheme Overview.
Figure 5. Online A/B test experimental results of DAU and LT.
Table 1. The performance of Online A/B Testing at the platform. Boldface represents significance level $p$-value <0.05<0.05< 0.05 of comparing KLAN with the online baseline. The values presented in the table represent the average observations of the last 7 days.
@@ 5.2. Online A/B Test
5.3. Online Experimental Results
5.3.1. Online A/B test results
(15) $\text{LT}=\frac{\sum_{i=\max(T-6,T_{0})}^{T}\text{DAU}_{i}}{\text{Total number of active users from }T_{0}\text{ to }T}$
Figure 5 illustrates the temporal evolution of DAU and LT growth rates throughout the 74-day experimental period, comprising a 14-day AA test followed by a 60-day AB test. During the AB test, we can find that KLAN achieves significant improvements in both metrics. Table 1 reports statistically significant improvements across all metrics compared to the online baseline. Notably, APP usage time, watch time and video view counts show substantial growth, while page drop-off ratio decreases significantly, indicating enhanced user satisfaction with assigned landing pages. Regarding platform constraint metrics, the effective page entry frequency metrics for all three page (Featured, Follow, and Explore) show positive improvements. Latency results indicate our model maintains real-time efficiency with negligible overhead. Currently, KLAN is fully deployed on Kuaishou, serving hundreds of millions of daily active users every day, demonstrating both its effectiveness and scalability in real-world applications.
Table 2. The ablation result of Online A/B Testing at the platform. Boldface represents significance level $p$-value <0.05<0.05< 0.05 of comparing KLAN variants with the online baseline.
@@ 5.3.2. Ablation Test of KLAN
Both individual modules outperform the baseline, highlighting the value of modeling multi-temporal (inter-day and intra-day) user signals. Notably, KLAN-ISP captures users’ inter-day static preferences, providing a strong foundation for navigation, while KLAN-IIT models intra-day dynamics, yielding consistent gains. Most importantly, the full KLAN model significantly surpasses both variants, validating our hierarchical design: ISP offers stable preferences, IIT captures short-term transitions, and their adaptive fusion via the AM module enables synergistic performance beyond individual contributions. These consistent improvements confirm KLAN’s effectiveness in balancing long-term and short-term user interests for optimal page navigation.

### v2

@@ Abstract.
Personalized Landing Page; Uplift Model; Reinforcement Learning
journalyear: 2026copyright: ccconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 09–13, 2026; Jeju Island, Republic of Koreabooktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’26), August 09–13, 2026, Jeju Island, Republic of Koreadoi: 10.1145/3770854.3783938isbn: 979-8-4007-2258-5/2026/08ccs: Information systems Recommender systems
@@ 1. Introduction
With the prevalence of Web 2.0 and mobile devices, a massive number of users have joined online platforms such as Amazon, TikTok, and Kuaishou for shopping, content creation, and social networking. On these platforms, multiple functional and channel pages are configured to cater to users’ diverse needs. For example, Kuaishou offers various types of pages including the ”Featured” page that presents recent and personalized recommendations to users and the ”Explore” page that provides less personalized but more diverse video recommendations. Inside each of these pages, users interact with the recommender system with consecutive requests, receiving sequential recommendations and providing feedback (Cui et al., 2025; Tang et al., 2023; Wu et al., 2025; Zhai et al., 2025b, a). In this sense, we consider the multi-page online recommender system as a two-stage interaction paradigm between the user and the platform (shown in Figure 1): an initial page navigation stage that navigates users to a specific functional or channel page, and a subsequent in-page interaction stage where users sequentially engage with customized content on the specific page.
@@
While most existing approaches for video recommendation platforms focus on the sequential recommendation task in the second in-page interaction stage (Kang and McAuley, 2018; Sun et al., 2019; Mao et al., 2025, 2024; Guo et al., 2023; Meng et al., 2023a, b, 2024; Chang et al., 2023; Pi et al., 2020; Li et al., 2024, 2025; Yu et al., 2025), there has been little investigation on how to achieve better page navigation in the first stage. However, the initial page navigation stage significantly impacts the user experience in the system. Firstly, it determines whether the following interactions can occur and directly influences the user’s first impression of the online service, which is reflected by short-term metrics like the Page Drop-off Ratio (PDR), as well as the long-term metrics such as Daily Active Users (DAU) and LifeTime (LT). Here, we define the Page Drop-off Ratio (PDR) as the probability that a user immediately leaves the assigned initial page upon app entry. Furthermore, users may typically enter the app multiple times in one day (Figure 2-c), so effective page navigation may choose different initial pages that accurately address the users’ dynamic demands upon each entry. Yet, inaccurate navigation could cause the users to frequently switch pages (i.e. the page drop-off), so it is also critical to avoid this negative impact on the user experience. Additionally, empirical evidence from Kuaishou, as shown in Figure 2-a and 2-b, demonstrates that approximately 42% of users engage across multiple pages and consistently outperform single-page users in terms of DAU and LT. While the remaining 58% of users exhibit single-page activities within a single day (shown in Figure 2-b), we believe that when the system properly introduces new pages to these users, it is possible to stimulate their potential in multi-page consumption, which improves the exposure of diverse video contents on different pages and eventually motivates the sustainable advancement of the platform’s ecosystem.
@@
  • Inter-day static page preferences modeling: User’s static page preference expresses the long-term stable preferences towards each pages, which is a critical feature to model the landing page of a user’s first entry on each day. Empirically, users have imbalanced interactions across pages, which introduces considerable exposure bias. For example, we observe that the vast majority of samples are available on the ”Featured” page. On the other hand, different pages are heterogeneous in presentation format, including content and interaction patterns, which results in vastly different user response distributions across pages (Section A.1). Given such heterogeneity, modeling user engagement for each page in isolation fails to capture the underlying causal relationships. Instead, a holistic perspective is needed—one that explicitly accounts for how assigning a user to one page over another influences their subsequent behavior. While uplift modeling provides a principled framework for estimating such treatment effects, existing approaches often rely on joint optimization over multiple objectives, which introduces gradient conflict problem, ultimately leading to suboptimal performance in practice.
@@ 2.1. Uplift Model for Video Recommendation Platform
Uplift models aim to identify the target user groups for each specific treatment by estimating individual treatment effects (ITE). Existing approaches can be categorized into three main methodologies: Meta-learner based methods, Tree-based methods and Neural Networks based methods. Meta-learner based methods (Künzel et al., 2019; Yao et al., 2021) constitute a classical framework for uplift modeling. The S-learner (Künzel et al., 2019) treats the treatment variable as a feature within a single model, while the T-learner (Künzel et al., 2019) builds separate models for treatment and control groups. Tree-based methods (Athey and Imbens, 2016; Nandy et al., 2023; Radcliffe and Surry, 2011; Wang et al., 2015) utilize specific tree or forest structures, setting direct computational methods for incremental uplift modeling. Due to the ability to capture complex treatment-response relationships and estimate more accurate ITE, neural networks have emerged as a popular framework for uplift modeling. Existing Neural Network-based uplift methods (Bica et al., 2020; Louizos et al., 2017; Shi et al., 2019; Wei et al., 2024; Zhou et al., 2023; Ke et al., 2021; Zhong et al., 2022) can be categorized into two categories: joint modeling of treatment and modeling of specific treatment causal effects. The former estimates the non-treatment score through features and incorporates the treatment feature into the treatment modeling based on this score. For instance, EFIN (Liu et al., 2023) captures the feature subset related to the control group through explicit feature selfinteraction, then models the sensitivity of non-treatment features to specific treatments using a treatment-aware interaction module, and enhances the model’s robustness through an intervention constraint module. The latter focuses on modeling the causal effects of specific treatment. CDUM (Meng et al., 2024) expands the treatment representation for guidance and indicator within a multi-task framework and proposes an online real-time feature learning network based on video recommendation scenarios. However, existing methods tend to either overemphasize the collaborative effects or the causal effects, making it difficult to adapt to short video recommendation scenarios which involves heterogeneous multi-treatment.
@@ 2.2. Reinforcement Learning for Video Recommendation Platform
Unlike traditional methods optimizing for the one-step recommendation, reinforcement learning (Sutton and Barto, 1998) approaches model video recommendation as a Markov Decision Process, optimizing cumulative rewards to maximize long-term user satisfaction. Reinforcement learning (RL) approaches can be broadly categorized into on-policy (Rummery and Niranjan, 1994; Mnih et al., 2016) and off-policy (Haarnoja et al., 2018; Watkins, 1989; van Hasselt et al., 2015) methods. Due to the high policy variance and low sample efficiency of on-policy methods, off-policy methods have emerged as the predominant choice in industrial application. Among off-policy methods, policy gradient approaches (Chen et al., 2019a, b; Ge et al., 2021, 2022; Li et al., 2022; Xian et al., 2019) suffer from the mismatch between the behavior policy and the target policy, leading to substantial discrepancies between offline evaluation and online performance. Value gradient approaches (van Hasselt et al., 2015; Pei et al., 2019; Taghipour et al., 2007; Zhao et al., 2021, 2018), prominently represented by DQN (van Hasselt et al., 2015), have demonstrated superior stability and data efficiency by leveraging experience replay. However, the traditional DQN struggles with the overestimation bias due to the exposure imbalance in our proposed PLPM scenario. To address it, conservative reinforcement learning algorithms such as CQL (Kumar et al., 2020) and IQL (Kostrikov et al., 2021) incorporate pessimistic value estimation to handle distributional shift. Despite the theoretical advantages of CQL, its practical deployment reveals a critical hyperparameter sensitivity issue. Specifically, an excessively large conservative coefficient leads to overly pessimistic policies that fail to stimulate multi-page user engagement, while an insufficient coefficient approximates standard DQN behavior, resulting in policy instability that violates online serving constraints and degrades user experience. Furthermore, our empirical analysis on Kuaishou reveals that intra-day RL decision exhibit a ”tidal phenomenon” (detailed in Appendix A.4). Motivated by this, we introduce a Dynamic Conservative Coefficient, leveraging intra-day real-time features strongly correlated with the ”tidal phenomenon” (such as request hour and daily app entry frequency) to automatically balance the exploration-exploitation trade-off.
@@ 3. The PLPM Problem Formulation
We define the three sub-tasks of PLPM problem as the inter-day modeling of user’s static page preferences, the intra-day modeling of user’s dynamic page interest transition and the adaptive modulation of both components.
@@ 3.2. The Intra-day Modeling Task
  • $\mathcal{S}$: The continuous representation space of the user state. Each state $s_{t}$ is composed of the real-time contextual features and the historical statistical features at step $t$.
@@
The goal of this MDP is to learn an optimal policy $\pi^{*}$$\mathcal{S}\rightarrow\mathcal{A}$ that maximizes the expected discounted return:
@@ 4.1.2. Multi-treatment Joint Optimization
Through Treatment-Specific Uplift Model, we obtain the estimated output pairs $\hat{y}_{i}^{k},\hat{y}_{i}^{0,k}]$ and the latent representation pairs $\mathbf{z}_{i}^{k},\mathbf{z}_{i}^{0,k}]$ from each branch. For training, we adopt the mean square error as the main training loss. To improve the robustness and effectiveness of ISP, we apply KL divergence to constrain non-treatment representations across branches. Therefore, the total training loss for ISP is as follows:
(5) $\mathcal{L}=\begin{cases}\frac{1}{N}\sum_{i=1}^{N}\sum_{k^{\prime}=1}^{K}\left(\left(y_{i}^{k}-\hat{y}_{i}^{0,k^{\prime}}\right)^{2}+D_{\mathrm{KL}}\left(z_{i}^{0,k^{\prime}}\,\middle\|\,{z}_{i}^{0,*}\right)\right),&k=0\\ \frac{1}{N}\sum_{i=1}^{N}\left(y_{i}^{k}-\hat{y}_{i}^{k}\right)^{2},&k\neq 0\end{cases}$
@@ 4.1.3. Prediction
During inference, we iteratively replace the treatment feature $\mathbf{t}_{i}$ to obtain predictions for all treatments: $\{\hat{y}_{i}^{k^{\prime}},\hat{y}_{i}^{0,k^{\prime}}\}_{k^{\prime}=0}^{K}$.
@@ 5.1. Dataset
Figure 4. Illustration of the online pipeline.
Figure 5. Online A/B test experimental results of DAU and LT.
@@
To model users’ intra-day dynamic interest transitions in the PLPM scenario, we construct an hourly-level dataset tailored for reinforcement learning (RL) tasks. This dataset is used to train the module KLAN-IIT, which captures temporal patterns in user behavior across app entry sessions. Each instance corresponds to a complete ”app entry session” (defined as the sequence of interactions that starts with the display of a landing page when the user enters the app and ends when the user exits) and the required features of the corresponding next session. The dataset is organized in the standard RL format $(s_{t},a_{t},r(s_{t},a_{t}),s_{t+1})$, where $a_{t}$ represents the assignment of a specific landing page upon app entry. The state $s_{t}$ includes both real-time contextual features—such as current-day page-level consumption statistics, the number of live broadcasts by followed creators at the time of entry, and the number of prior app entries on the same day—and historical statistical features, including aggregated page consumption behavior over the past $n$ days. The reward $r(s_{t},a_{t})$ is defined as the app usage duration of the current session, serving as a proxy for user engagement. This dataset is constructed from hourly online logs collected in a real-world industrial environment, ensuring high temporal resolution and behavioral fidelity.
@@
Table 1. The performance of Online A/B Testing at the platform. Boldface represents significance level $p$-value <0.05<0.05 of comparing KLAN with the online baseline. The values presented in the table represent the average observations of the last 7 days.
@@
5.2. Online A/B Test
5.2.1. System Description.
Table 2. The ablation result of Online A/B Testing at the platform. Boldface represents significance level $p$-value <0.05<0.05 of comparing KLAN variants with the online baseline.
@@ 5.2. Online A/B Test
5.2.2. Scheme Overview.
5.3. Online Experimental Results
5.3.1. Online A/B test results
[…]
Figure 5 illustrates the temporal evolution of DAU and LT growth rates throughout the 74-day experimental period, comprising a 14-day AA test followed by a 60-day AB test. During the AB test, we can find that KLAN achieves significant improvements in both metrics. Table 1 reports statistically significant improvements across all metrics compared to the online baseline. Notably, APP usage time, watch time and video view counts show substantial growth, while page drop-off ratio decreases significantly, indicating enhanced user satisfaction with assigned landing pages. Regarding platform constraint metrics, the effective page entry frequency metrics for all three pages show positive improvements. Latency results indicate our model maintains real-time efficiency with negligible overhead. Currently, KLAN is fully deployed on Kuaishou, serving hundreds of millions of daily active users every day.
@@ 5.3.2. Ablation Test of KLAN
Both individual modules outperform the baseline, highlighting the value of modeling multi-temporal (inter-day and intra-day) user signals. Notably, KLAN-ISP captures users’ inter-day static preferences, providing a strong foundation for navigation, while KLAN-IIT models intra-day dynamics, yielding consistent gains. Most importantly, KLAN significantly surpasses both variants, validating our hierarchical design: ISP offers stable preferences, IIT captures short-term transitions, and their adaptive fusion via the AM module enables synergistic performance beyond individual contributions. These consistent improvements confirm KLAN’s effectiveness in balancing long-term and short-term user interests for optimal page navigation.